OpenStudy (anonymous):

how do you factor this equation? x^3+6x^2+12x+8 = 0

OpenStudy (amistre64):

with a ten foot pole ?

OpenStudy (amistre64):

start by getting number that might work by factoring the last over the first 1,8,2,4 ------ makes life a little easier 1

OpenStudy (anonymous):

factor by grouping

OpenStudy (amistre64):

since all the terms are + already we know there are no positive solutions to be had so lets go with - tests

OpenStudy (amistre64):

lets try x = -2 -2 | 1 6 12 8 0 -2 -8 -8 ----------- 1 4 4 0 <- its a good one casue we got 0 i n the end

OpenStudy (amistre64):

x=-2 is a solution; (x+2) (x^2+4x+4) the quad factors quite nicely as well to get: (x+2)(x+2)(x+2) = (x+2)^3

OpenStudy (anonymous):

why did you use -2? not 2

OpenStudy (anonymous):

also how did you get (x+2)

OpenStudy (anonymous):

or: (x^3+6x^2)+(12x+8)=0 x^2(x+6)+4(3x+2)=0 (x^2+4)(x+6)(3x+2)=0 x=-6 x=-2/3 x^2=4 \[x=\sqrt{4} = \pm 2\] then plug check to see which one solution, I think x=-2 is solution

OpenStudy (anonymous):

Algebra I style : )

OpenStudy (anonymous):

thank you for your medal

OpenStudy (amistre64):

the rule for wigns tells me that if all are + then there are no poistive roots to be had; so I went with the negative versions of my gene pool

OpenStudy (amistre64):

rule for signs.... its like im having a stroke lol

OpenStudy (anonymous):

have to set factors equal to zero and solve...so much easier doing this stuff than my own...grrhhhh, lol!

OpenStudy (anonymous):

the way amis do is call long division,you can do that way too ,but is hard to explant

OpenStudy (amistre64):

x = sqrt(4) = 2 NOT +-2

OpenStudy (amistre64):

not long division; close tho; synthetic division ;)

OpenStudy (anonymous):

square= +-

OpenStudy (amistre64):

sqrt(4) = 2 2^2 = +- sqrt(4)

OpenStudy (amistre64):

got it lol i read it wrong

OpenStudy (anonymous):

yes ,almos close

OpenStudy (amistre64):

now the question asked to factor it; not find roots right?

OpenStudy (amistre64):

(x+2)^3 is the factored form then